OpenStudy (anonymous):

If u want to see how far something slides in physics what formula can I use. With the given velocity initial and velocity final and time

OpenStudy (anonymous):

how far something slides in physics?

OpenStudy (amistre64):

well, you can determine the acceleration of the system and from that develop a position equation

OpenStudy (amistre64):

acceleration is the change in velocity over a given time interval. from there its just working up a quadratic with integrations

OpenStudy (anonymous):

How far something is? Distance

OpenStudy (amistre64):

*m/s - *m/s ------------- = k m/s^2 *s if a(t) = k v(t) = kt + c p(t) = kt^2/2 + ct + d

OpenStudy (anonymous):

Thanks that helped allot

OpenStudy (amistre64):

youre welcome